# ULog: Unsupervised Log Parsing with Large Language Models through Log Contrastive Units

### v1

@@ Abstract.
Log parsing serves as an essential prerequisite for various log analysis tasks. Recent advancements in this field have improved parsing accuracy by leveraging the semantics in logs through fine-tuning large language models (LLMs) or learning from in-context demonstrations. However, these methods heavily depend on labeled examples to achieve optimal performance. In practice, collecting sufficient labeled data is challenging due to the large scale and continuous evolution of logs, leading to performance degradation of existing log parsers after deployment. To address this issue, we propose ULog, an unsupervised LLM-based method for efficient and off-the-shelf log parsing. Our key insight is that while LLMs may struggle with direct log parsing, their performance can be significantly enhanced through comparative analysis across multiple logs that differ only in their parameter parts. We refer to such groups of logs as Log Contrastive Units (LCUs). Given the vast volume of logs, obtaining LCUs is difficult. Therefore, ULog introduces a hybrid ranking scheme to effectively search for LCUs by jointly considering the commonality and variability among logs. Additionally, ULog crafts a novel parsing prompt for LLMs to identify contrastive patterns and extract meaningful log structures from LCUs. Experiments on large-scale public datasets demonstrate that ULog significantly outperforms state-of-the-art log parsers in terms of accuracy and efficiency, providing an effective and scalable solution for real-world deployment.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
Log messages record events, transactions, or activities generated by software applications and operating systems at runtime (Wang et al., 2022; He et al., 2022b; Liu et al., 2023a). They provide valuable insights for system performance monitoring and reliability assurance. Various tools have been developed to conduct automated log analysis tasks, including anomaly detection (Zhang et al., 2019; Zhao et al., 2021; Zhang et al., 2022; Ali et al., 2023; Chen et al., 2021b), root cause analysis (Amar and Rigby, 2019; Wang et al., 2020; Notaro et al., 2023), and failure diagnosis (Xu et al., 2009; Chen et al., 2021a). Log parsing, which transforms semi-structured log messages into structured formats (Khan et al., 2022), serves as a critical preliminary step in log analysis. Typically, a raw log message contains two parts: 1) log templates: constant parts that describe the main content of the logged event; 2) log parameters: dynamic parts that contain the parameters (determined at runtime) associated with the event. Figure 1 demonstrates some log messages generated by their logging statements and parsed into structured data.
@@
To address this label-demanding problem, we propose ULog, an unsupervised LLM-based log parser, which can generalize to any new log dataset without manual annotations or prior knowledge of the specific log formats. The core idea behind LUNAR lies in leveraging LLMs’ ability to perform comparative analysis on multiple log messages that vary in their parameter parts. While LLMs may struggle with direct log parsing, the comparison can derive valuable insights by identifying and interpreting patterns across these variations. For instance, by contrasting logs such as “session opened for user news” and “session opened for user test,” LLMs can easily infer that the tokens “news” and “test” are likely to represent a username parameter. We refer to such group of logs as a Log Contrastive Unit (LCU), which are similar enough to facilitate meaningful comparisons, yet diverse enough to highlight the variable parameters. To find LCUs, we introduce a hybrid ranking scheme by jointly considering the commonality and variability among logs. However, given the vast volume of logs, evaluating every possible combination of logs is impractical. Thus, we introduce a hierarchical sharder to first divide logs into different buckets based on log length and top-$k$ frequent tokens. The LCU selection and subsequent log parsing can then be efficiently performed in each bucket in parallel. With the selected LCU, ULog crafts a novel specialized parsing prompt without the need of labelled templates for ICL. The prompt specifies task intention and output constraints in detail, and provides representative parameter examples to inform LLMs of the parameter characteristics.
We evaluate the performance of ULog against a range of label-free (unsupervised) parsers and label-dependent parsers. The experiments are conducted on 14 large-scale log parsing datasets in Loghub-2.0 (Jiang et al., 2023b) from LogPAI (Zhu et al., 2019). The results show that ULog substantially outperforms the unsupervised baselines, surpassing Brain (Yu et al., 2023) and LILAC w/o ICL (Jiang et al., 2023a) by 46.2% and 26.9% in FTA, respectively. When compared with label-dependent methods, ULog achieves performance on par with the current state-of-the-art parser, LILAC. Moreover, with parallelization capability, ULog attains a parsing speed comparable to most syntax-based baselines and superior than semantic-based baselines, facilitating efficient parsing of large-scale log data. Our evaluation shows the potential of ULog for deployment in real-world production systems, where the efficiency, accuracy and generalizability are critical concerns.
@@
  • To the best of our knowledge, we propose the first unsupervised LLM-based log parser dubbed ULog, which instructs LLMs to make comparisons on log contrastive units (LCUs).
@@
  • We evaluate ULog on large-scale public datasets. The results show that ULog significantly outperforms unsupervised parsers and achieves comparable performance with state-of-the-art LILAC in terms of accuracy and efficiency.
@@ 3. Methodology
Figure 4. The overall workflow of ULog.
@@ 3.1. Overview
Figure 4 illustrates the overall framework of ULog, which consists of three main components: hierarchical sharder, generative LCU ranker, and label-free LLM parser. To address the first challenge, we propose the hierarchical sharder, which divides raw log messages into different buckets based on log length and top-$k$ ranked tokens. By separating logs with low similarity into different buckets, this component reduces the sampling overhead and enables parallelization for efficient parsing (§3.2). To address the second challenge, we propose the generative LCU ranker, which operates within each bucket to continuously sample LCUs for the LLM to parse. This module computes a hybrid LCU score that jointly considers both the commonality and variability of the LCUs, guiding the sampling process to ensure effective LCUs (§3.3). Lastly, the label-free LLM parser constructs an organized prompt for the mined LCUs to query the LLM and obtain the templates from the LLM’s response. We introduce a novel format for organizing the parsing prompt, which specifies the task intention and output constraints, and includes several representative parameter examples to inform the LLM of the parameter characteristics (§3.4). The combination of all components ensures that the LLM can effectively parse the logs without the need for labelled data, addressing both scalability and diversity challenges.
Figure 5. Hierarchical log sharding in ULog.
@@ 3.2. Hierarchical Log Sharder
Directly sampling LCUs from the whole log messages is overwhelming due to the large volume of log data, which often contain million of log messages. Hence, ULog first groups the log messages into multiple log buckets, where each bucket contain logs that share some similarity. By doing so, extremely different logs, which have less potentials to belong to the same template (He et al., 2017), are placed to different buckets. Specifically, we propose a hierarchical algorithm to shard log: 1) firstly divide the log messages by log length, and 2) progressively divide logs by top-$k$ frequent tokens. Figure. 5 demonstrates the workflow of hierarchical log sharder.
@@
ULog first shards the log messages by the log length. Logs with the same number of tokens are grouped in a bucket, which is a widely adopted heuristic to conduct initial grouping in log analysis (He et al., 2017; Li et al., 2023b). Here we simply adopt the white-space as the delimiter. We do not use other delimiters such as ‘:‘ because most of these delimiters appear in parameters and introducing them can produce over-fragmented logs, leading to a noisy sharding result.
@@
However, simply sharding by log length is too coarse-grained since logs belonging to different templates could have the same length (Jiang et al., 2023a). Thus ULog continuously shards each bucket into more purified ones where less templates are included in one bucket. Intuitively, the static parts of a log generally have higher occurrences than its parameter part, therefore logs that share the most frequent tokens have more potential to belong to the same template (Liu et al., 2019b; Jiang et al., 2023b, a). To achieve this, we introduce a hierarchical agglomerative clustering algorithm based on top-$k$ frequent tokens, which builds the sub-buckets in a bottom-up way.
@@ 3.3. Two-stage LCU Selector
After log sharding, ULog iteratively select a group of similar logs for the LLM to make comparisons and extract templates. A suitable group of logs should be similar in tokens to each other, as logs generated by the same logging statement share several identical tokens. On this basis, the logs are expected to have as much variance as possible, enabling LLMs to make cross-log comparisons to infer parameters. For example, in Figure 6, group ➀ and ➁ are preferred than group ➂ as their logs are more likely to belong to the same template. Moreover, group ➀ is better than ➁ due to its higher variance reflected in the diverse parameter value of process index.
@@
Given a bucket of logs, ULog first generate multiple candidate LCUs that share some similarity. A straightforward approach is to enumerate all possible log combinations. However, this method is computationally expensive due to the sheer volume of logs and the complexity of enumeration. To reduce computation overhead, we apply stratified sampling to collect a limited-sized log pool by sampling from different similarity levels before perform combination.
Given a log bucket, ULog first randomly selects a log message in the bucket as the anchor log. Next, ULog computes the pair-wise similarities of the anchor log to the remaining logs. We use Jaccard similarity, which is a widely adopted metric to measure pair-wise log similarity (He et al., 2017; Jiang et al., 2023a). Specifically, we first split the log $l$ into tokens $T(l)$ with white-space delimiter and then compute the Jaccard similarity $JS(l_{1},l_{2})=\frac{T(l_{1})\cap T(l_{2})}{T(l_{1})\cup T(l_{2})}$. Then, we remove the logs with a similarity below a threshold $s$ or equals to $1.0$. The poor similarity indicates the less possibility to belong to the same template, while a similarity of $1.0$ indicates duplication. After that, we create a pool of logs by stratified sampling from each similarity level to ensure balanced sampling. Specifically, we set the sample size of each level equal to the LCU size $m$ to involve more diverse logs. Lastly, from the log pool, ULog generates multiple LCUs by combinatorial enumeration. Each LCU consists of an anchor log, as well as $m-1$ logs chosen from the log pool. Since we sample a small size of LCU (e.g., $m=3$ or $m=4$) from a limited size of log pool, the computational cost of combinatorial enumeration is within an acceptable range (More details about efficiency can found in Section 5.3). If the log pool contain less than $m-1$ logs, we directly return the anchor with the pool as the candidate LCU. Finally, a candidate set of LCUs are generated and proceeded to hybrid rankding in the next step.
@@ 3.3.2. Hybrid LCU Ranking
After obtaining a small number candidate LCUs, ULog conducts hybrid ranking to select the most suitable LCU. The selected logs are required to exhibit variability, while maintain some commonality as we discussed in Section 2.2. By doing so, LLMs can be less likely to incorrectly recognize the frequent tokens as parameters, while also benefit from the contrastive information in the LCU. Specifically, we propose to compute the variability score and the commonality score to measure an LCU.
@@ 3.4. Label-free LLM Parser
Upon obtaining an LCU, ULog applies an LLM-based parser to extract the template without any labelled examples. The parser leverages the strong textual comprehension ability of LLMs, which have shown remarkable few-shot and zero-shot performance in various information extraction tasks, such as named entity recognition (Xie et al., 2023) and text classification (Lyu et al., 2023). Therefore, we believe LLMs have the potential to solve log parsing in an unsupervised way. Specifically, our LLM parser first creates a well-designed prompt for each LCU to instruct the LLM (§3.4.1) and then extract the template from the LLM response (§3.4.2).
@@ 3.4.1. Prompt Design
Figure 7. Demonstration-free parsing prompt in ULog.
@@
Parameter Examples. Inspired by the entity examples used for unsupervised NER (Jiang et al., 2024), we involve parameter examples to the prompt. These examples are used to provide domain knowledge and instruct the task, which consist of a parameter type and a parameter value. For instance, “/var/www/html/” is the value of a parameter “directory.” To obtain the parameter examples, we first manually check the regular expressions of used in previous studies (He et al., 2017; Li et al., 2023a), then sample a value from matched parameters, and finally summarize the parameter type of the value (Li et al., 2023a). In the prompt, the parameter value and type are connected by an arrow, aiming to indicate the task intent to convert the value to a placeholder with the bracket.
@@ 3.5. Bucket Updating
After obtaining a valid template, ULog examines the logs that match the template. The template is assigned as the final prediction to these successfully matched logs, which are then removed from the buckets. The remaining log buckets are used for next iteration of parsing. Once all buckets are empty, the parsing terminates.
It is worth-noting that the the hierarchical sharding produces buckets at two levels, i.e., sharded by length and further clustered by top-$k$ tokens. A straightforward approach for bucket updating is to only examine and update the final bucket. However, logs in the same template can be placed in different buckets of equal length, potentially causing redundant iterations for one template, leading to prolonged parsing time and increased LLM querying expenses. To balance the parsing overhead, ULog conducts template examination and updating on the buckets under the first sharding level, meaning that buckets with identical log lengths are all examined once a template is obtained.

### v2

# LUNAR: Unsupervised LLM-based Log Parsing

@@ Abstract.
Log parsing serves as an essential prerequisite for various log analysis tasks. Recent advancements in this field have improved parsing accuracy by leveraging the semantics in logs through fine-tuning large language models (LLMs) or learning from in-context demonstrations. However, these methods heavily depend on labeled examples to achieve optimal performance. In practice, collecting sufficient labeled data is challenging due to the large scale and continuous evolution of logs, leading to performance degradation of existing log parsers after deployment. To address this issue, we propose LUNAR, an unsupervised LLM-based method for efficient and off-the-shelf log parsing. Our key insight is that while LLMs may struggle with direct log parsing, their performance can be significantly enhanced through comparative analysis across multiple logs that differ only in their parameter parts. We refer to such groups of logs as Log Contrastive Units (LCUs). Given the vast volume of logs, obtaining LCUs is difficult. Therefore, LUNAR introduces a hybrid ranking scheme to effectively search for LCUs by jointly considering the commonality and variability among logs. Additionally, LUNAR crafts a novel parsing prompt for LLMs to identify contrastive patterns and extract meaningful log structures from LCUs. Experiments on large-scale public datasets demonstrate that LUNAR significantly outperforms state-of-the-art log parsers in terms of accuracy and efficiency, providing an effective and scalable solution for real-world deployment. [^1] The code and data are available at https://github.com/Jun-jie-Huang/LUNAR.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; October 27 - November 1, 2024; Sacramento, US
@@ 1. Introduction
Log messages record events, transactions, or activities generated by software applications and operating systems at runtime (He et al., 2022b; Li et al., 2023a, 2024). They provide valuable insights for system performance monitoring and reliability assurance. Various tools have been developed to conduct automated log analysis tasks, including anomaly detection (Zhang et al., 2019; Zhao et al., 2021; Zhang et al., 2022; Liu et al., 2023a; Chen et al., 2021b), root cause analysis (Amar and Rigby, 2019; Wang et al., 2020; Notaro et al., 2023), and failure diagnosis (Xu et al., 2009; Chen et al., 2021a). Log parsing, which transforms semi-structured log messages into structured formats (Khan et al., 2022), serves as a critical preliminary step in log analysis. Typically, a raw log message contains two parts: 1) log templates: constant parts that describe the main content of the logged event; 2) log parameters: dynamic parts that contain the parameters (determined at runtime) associated with the event. Figure 1 demonstrates some log messages generated by their logging statements and parsed into structured data.
@@
To address this label-demanding problem, we propose LUNAR, an LLM-based unsupervised log parser, which can generalize to any new log dataset without manual annotations or prior knowledge of the specific log formats. The core idea behind LUNAR lies in leveraging LLMs’ ability to perform comparative analysis on multiple log messages that vary in their parameter parts. While LLMs may struggle with direct log parsing, the comparison can derive valuable insights by identifying and interpreting patterns across these variations. For instance, by contrasting logs such as “session opened for user news” and “session opened for user test,” LLMs can easily infer that the tokens “news” and “test” are likely to represent a username parameter. We refer to such group of logs as a Log Contrastive Unit (LCU), which are similar enough to facilitate meaningful comparisons, yet diverse enough to highlight the variable parameters. To find LCUs, we introduce a hybrid ranking scheme by jointly considering the commonality and variability among logs. However, given the vast volume of logs, evaluating every possible combination of logs is impractical. Thus, we introduce a hierarchical sharder to first divide logs into different buckets based on log length and top-$k$ frequent tokens. The LCU selection and subsequent log parsing can then be efficiently performed in each bucket in parallel. With the selected LCU, LUNAR crafts a novel specialized parsing prompt without the need of labelled templates for ICL. The prompt specifies task intention and output constraints in detail, and provides representative parameter examples to inform LLMs of the parameter characteristics.
We evaluate the performance of LUNAR against a range of label-free (unsupervised) parsers and label-dependent parsers. The experiments are conducted on 14 large-scale log parsing datasets in Loghub-2.0 (Jiang et al., 2023b) from LogPAI (Zhu et al., 2019). The results show that LUNAR substantially outperforms the unsupervised baselines, surpassing Brain (Yu et al., 2023) and LILAC w/o ICL (Jiang et al., 2023a) by 46.2% and 26.9% in FTA, respectively. When compared with label-dependent methods, LUNAR achieves performance on par with the current state-of-the-art parser, LILAC. Moreover, with parallelization capability, LUNAR attains a parsing speed comparable to most syntax-based baselines and superior than semantic-based baselines, facilitating efficient parsing of large-scale log data. Our evaluation shows the potential of LUNAR for deployment in real-world production systems, where the efficiency, accuracy and generalizability are critical concerns.
@@
  • To the best of our knowledge, we propose the first LLM-based unsupervised log parser dubbed LUNAR, which instructs LLMs to make comparisons on log contrastive units (LCUs).
@@
  • We evaluate LUNAR on large-scale public datasets. The results show that LUNAR significantly outperforms unsupervised parsers and achieves comparable performance with state-of-the-art LILAC in terms of accuracy and efficiency.
@@ 3. Methodology
Figure 4. The overall workflow of LUNAR.
@@ 3.1. Overview
Figure 4 illustrates the overall framework of LUNAR, which consists of three main components: hierarchical sharder, generative LCU ranker, and label-free LLM parser. To address the first challenge, we propose the hierarchical sharder, which divides raw log messages into different buckets based on log length and top-$k$ ranked tokens. By separating logs with low similarity into different buckets, this component reduces the sampling overhead and enables parallelization for efficient parsing (§3.2). To address the second challenge, we propose the generative LCU ranker, which operates within each bucket to continuously sample LCUs for the LLM to parse. This module computes a hybrid LCU score that jointly considers both the commonality and variability of the LCUs, guiding the sampling process to ensure effective LCUs (§3.3). Lastly, the label-free LLM parser constructs an organized prompt for the mined LCUs to query the LLM and obtain the templates from the LLM’s response. We introduce a novel format for organizing the parsing prompt, which specifies the task intention and output constraints, and includes several representative parameter examples to inform the LLM of the parameter characteristics (§3.4). The combination of all components ensures that the LLM can effectively parse the logs without the need for labelled data, addressing both scalability and diversity challenges.
Figure 5. Hierarchical log sharding in LUNAR.
@@ 3.2. Hierarchical Log Sharder
Directly sampling LCUs from the whole log messages is overwhelming due to the large volume of log data, which often contain million of log messages. Hence, LUNAR first groups the log messages into multiple log buckets, where each bucket contain logs that share some similarity. By doing so, extremely different logs, which have less potentials to belong to the same template (He et al., 2017), are placed to different buckets. Specifically, we propose a hierarchical algorithm to shard log: 1) firstly divide the log messages by log length, and 2) progressively divide logs by top-$k$ frequent tokens. Figure. 5 demonstrates the workflow of hierarchical log sharder.
@@
LUNAR first shards the log messages by the log length. Logs with the same number of tokens are grouped in a bucket, which is a widely adopted heuristic to conduct initial grouping in log analysis (He et al., 2017; Li et al., 2023c). Here we simply adopt the white-space as the delimiter. We do not use other delimiters such as ‘:‘ because most of these delimiters appear in parameters and introducing them can produce over-fragmented logs, leading to a noisy sharding result.
@@
However, simply sharding by log length is too coarse-grained since logs belonging to different templates could have the same length (Jiang et al., 2023a). Thus LUNAR continuously shards each bucket into more purified ones where less templates are included in one bucket. Intuitively, the static parts of a log generally have higher occurrences than its parameter part, therefore logs that share the most frequent tokens have more potential to belong to the same template (Liu et al., 2019b; Jiang et al., 2023b, a). To achieve this, we introduce a hierarchical agglomerative clustering algorithm based on top-$k$ frequent tokens, which builds the sub-buckets in a bottom-up way.
@@ 3.3. Two-stage LCU Selector
After log sharding, LUNAR iteratively select a group of similar logs for the LLM to make comparisons and extract templates. A suitable group of logs should be similar in tokens to each other, as logs generated by the same logging statement share several identical tokens. On this basis, the logs are expected to have as much variance as possible, enabling LLMs to make cross-log comparisons to infer parameters. For example, in Figure 6, group ➀ and ➁ are preferred than group ➂ as their logs are more likely to belong to the same template. Moreover, group ➀ is better than ➁ due to its higher variance reflected in the diverse parameter value of process index.
@@
Given a bucket of logs, LUNAR first generate multiple candidate LCUs that share some similarity. A straightforward approach is to enumerate all possible log combinations. However, this method is computationally expensive due to the sheer volume of logs and the complexity of enumeration. To reduce computation overhead, we apply stratified sampling to collect a limited-sized log pool by sampling from different similarity levels before perform combination.
Given a log bucket, LUNAR first randomly selects a log message in the bucket as the anchor log. Next, LUNAR computes the pair-wise similarities of the anchor log to the remaining logs. We use Jaccard similarity, which is a widely adopted metric to measure pair-wise log similarity (He et al., 2017; Jiang et al., 2023a). Specifically, we first split the log $l$ into tokens $T(l)$ with white-space delimiter and then compute the Jaccard similarity $JS(l_{1},l_{2})=\frac{T(l_{1})\cap T(l_{2})}{T(l_{1})\cup T(l_{2})}$. Then, we remove the logs with a similarity below a threshold $s$ or equals to $1.0$. The poor similarity indicates the less possibility to belong to the same template, while a similarity of $1.0$ indicates duplication. After that, we create a pool of logs by stratified sampling from each similarity level to ensure balanced sampling. Specifically, we set the sample size of each level equal to the LCU size $m$ to involve more diverse logs. Lastly, from the log pool, LUNAR generates multiple LCUs by combinatorial enumeration. Each LCU consists of an anchor log, as well as $m-1$ logs chosen from the log pool. Since we sample a small size of LCU (e.g., $m=3$ or $m=4$) from a limited size of log pool, the computational cost of combinatorial enumeration is within an acceptable range (More details about efficiency can found in Section 5.3). If the log pool contain less than $m-1$ logs, we directly return the anchor with the pool as the candidate LCU. Finally, a candidate set of LCUs are generated and proceeded to hybrid rankding in the next step.
@@ 3.3.2. Hybrid LCU Ranking
After obtaining a small number candidate LCUs, LUNAR conducts hybrid ranking to select the most suitable LCU. The selected logs are required to exhibit variability, while maintain some commonality as we discussed in Section 2.2. By doing so, LLMs can be less likely to incorrectly recognize the frequent tokens as parameters, while also benefit from the contrastive information in the LCU. Specifically, we propose to compute the variability score and the commonality score to measure an LCU.
@@ 3.4. Label-free LLM Parser
Upon obtaining an LCU, LUNAR applies an LLM-based parser to extract the template without any labelled examples. The parser leverages the strong textual comprehension ability of LLMs, which have shown remarkable few-shot and zero-shot performance in various information extraction tasks, such as named entity recognition (Xie et al., 2023) and text classification (Lyu et al., 2023). Therefore, we believe LLMs have the potential to solve log parsing in an unsupervised way. Specifically, our LLM parser first creates a well-designed prompt for each LCU to instruct the LLM (§3.4.1) and then extract the template from the LLM response (§3.4.2).
@@ 3.4.1. Prompt Design
Figure 7. Demonstration-free parsing prompt in LUNAR.
@@
Parameter Examples. Inspired by the entity examples used for unsupervised NER (Jiang et al., 2024), we involve parameter examples to the prompt. These examples are used to provide domain knowledge and instruct the task, which consist of a parameter type and a parameter value. For instance, “/var/www/html/” is the value of a parameter “directory.” To obtain the parameter examples, we first manually check the regular expressions of used in previous studies (He et al., 2017; Li et al., 2023b), then sample a value from matched parameters, and finally summarize the parameter type of the value (Li et al., 2023b). In the prompt, the parameter value and type are connected by an arrow, aiming to indicate the task intent to convert the value to a placeholder with the bracket.
@@ 3.5. Bucket Updating
After obtaining a valid template, LUNAR examines the logs that match the template. The template is assigned as the final prediction to these successfully matched logs, which are then removed from the buckets. The remaining log buckets are used for next iteration of parsing. Once all buckets are empty, the parsing terminates.
It is worth-noting that the the hierarchical sharding produces buckets at two levels, i.e., sharded by length and further clustered by top-$k$ tokens. A straightforward approach for bucket updating is to only examine and update the final bucket. However, logs in the same template can be placed in different buckets of equal length, potentially causing redundant iterations for one template, leading to prolonged parsing time and increased LLM querying expenses. To balance the parsing overhead, LUNAR conducts template examination and updating on the buckets under the first sharding level, meaning that buckets with identical log lengths are all examined once a template is obtained.